OpenStudy (anonymous):

What is the area of a circle with a diameter of 8? A. 4π B. 8π c. 16π d. 64π

OpenStudy (mathstudent55):

Do you know the formula for the area of a circle?

OpenStudy (lsdxmdmaddicthc):

The formula for Area of a Circle is : A = (pi) * r(squared)

OpenStudy (lsdxmdmaddicthc):

r = radius, which is half of the diameter (which you already know)

OpenStudy (mathstudent55):

\(\large A_{circle} = \pi r^2\) where r = radius You were given the diameter. The radius is half of the diameter. Find the radius first, then use the formula.

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

@IrisCavadraVenda Direct answers aren't allowed, and A is incorrect.

OpenStudy (anonymous):

Sorry! I didnt know I havent been on open study in a long time I will remember that

OpenStudy (anonymous):

@floatingsugar Any ideas?

OpenStudy (anonymous):

my answeer is c

OpenStudy (anonymous):

@srebollozo Direct answers aren't allowed.

OpenStudy (anonymous):

So it would be C? That is what i got when I did 3.14*4 squared. It took me to 16, correct?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

so would be c? i got divide that 25.12 by 8 gonna be 16